OpenStudy (anonymous):

rx+9/5=h please help!!!!!!!! The division bar would be a fraction so it would be rx+9 for the numerator and 5 for the denominator, please help

OpenStudy (kayders1997):

Solve for?

OpenStudy (anonymous):

solve for x

OpenStudy (kayders1997):

So this was kinda like the first one you asked so you have to do the adding and subtracting first so what would you have to move to the other side to get x closer to being by itself?

OpenStudy (kayders1997):

Do you understand?

OpenStudy (kayders1997):

You would have to subtract 9/5 and bring it to the other side

OpenStudy (kayders1997):

Do you need me to draw it out?

OpenStudy (anonymous):

yes please

OpenStudy (kayders1997):

https://www.twiddla.com/2550825

OpenStudy (kayders1997):

Just click on the link

OpenStudy (kayders1997):

I can't add a drawing because I'm on an ipad

OpenStudy (kayders1997):

Or can't you?

OpenStudy (kayders1997):

I can't help you if you don't click on the link, it's only a website so you can see what I'm doing with the problem, it's not going to hurt your computer

OpenStudy (kayders1997):

I can't attach a file or do a drawing for some reason on here @j.nejad

OpenStudy (anonymous):

thank you

OpenStudy (kayders1997):

Don't want to use the link?

OpenStudy (kayders1997):

You don't have to sign up or anything

OpenStudy (kayders1997):

Than I can explain really fast too

OpenStudy (kayders1997):

Sorry

OpenStudy (anonymous):

could u go on the link?

OpenStudy (kayders1997):

I hope that helps i'm sorry i'm so tired

OpenStudy (kayders1997):

So you do you adding/subtracting first thab multiplacation/division next

OpenStudy (kkutie7):

\[\frac{rx+5}{9}=h\] you need to do multiplication or division first. In order to do addition or subtraction there needs to be more than one piece of information on the left hand side. please let me see your first step so I can help you.

OpenStudy (kkutie7):

I have to log off soon so if you want my help we need to hurry this up :P

OpenStudy (kkutie7):

*explanation examples* division or multiplication first: \[\frac{1x}{2}=y \rightarrow 1x=2y\rightarrow x=2y\] \[5x=y\rightarrow x=\frac{y}{5}\] Addition or subtraction: \[2x+5=y\rightarrow 2x=y-5 \rightarrow x=\frac{y-5}{2}\] or \[2x-5=y\rightarrow 2x=y+5 \rightarrow x=\frac{y+5}{2}\] A little bit more complicated. Addition and subtraction with fractions: \[\frac{7x}{2}+\frac{3}{5}=y\rightarrow \frac{7x}{2}=y-\frac{3}{5}\rightarrow 7x=2(y-\frac{3}{5})\rightarrow \frac{2(y-\frac{3}{5})}{7}=x\]